# Citekit: A Modular Toolkit for Large Language Model Citation Generation

### v1

@@ Abstract
Enabling Large Language Models (LLMs) to generate citations in Question-Answering (QA) tasks is an emerging paradigm aimed at enhancing the verifiability of their responses when LLMs are utilizing external references to generate an answer. However, there is currently no unified framework to standardize and fairly compare different citation generation methods, leading to difficulties in reproducing different methods and a comprehensive assessment. To cope with the problems above, we introduce Citekit , an open-source and modular toolkit designed to facilitate the implementation and evaluation of existing citation generation methods, while also fostering the development of new approaches to improve citation quality in LLM outputs. This tool is highly extensible, allowing users to utilize 4 main modules and 14 components to construct a pipeline, evaluating an existing method or innovative designs. Our experiments with two state-of-the-art LLMs and 11 citation generation baselines demonstrate varying strengths of different modules in answer accuracy and citation quality improvement, as well as the challenge of enhancing granularity. Based on our analysis of the effectiveness of components, we propose a new method, self-RAG Snippet, obtaining a balanced answer accuracy and citation quality. Citekit is released at https://github.com/SjJ1017/Citekit.
@@
Jiajun Shen1,3*, Tong Zhou1*, Suifeng Zhao4, Yubo Chen1,2, Kang Liu1,2,5 1The Key Laboratory of Cognition and Decision Intelligence for Complex Systems Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3University of Chinese Academy of Sciences 4Beijing University of Posts and Telecommunications 5Shanghai Artificial Intelligence Laboratory shenjiajun21@mails.ucas.ac.cn, tong.zhou@ia.ac.cn, zhaosuifeng@bupt.edu.cn, {yubo.chen, kliu}@nlpr.ia.ac.cn
@@ 1 Introduction
Given the urgent need, ALCE (Gao et al., 2023b) is the first attempt systematically to develop some basic methods to enable LLMs to generate citations in QA tasks and propose metrics for evaluating the quality of citations, showing there is still room for improvement concerning citation generation. Following ALCE’s contribution, there are other methods that either use training (Huang et al., 2024; Li et al., 2024a; Ye et al., 2024) or construct complicated pipelines to enhance the ability of generative models in citing external documents (Zhang et al., 2024; Hennigen et al., 2024; Lee et al., 2023). For example, Fierro et al. (2024) found the black-box generation is not factually faithful, so they use blueprint models to generate plans or blueprints and the output can be traced back to the blueprint. Verifiable Text Generation (VTG) (Sun et al., 2024) uses verifiers, evidence finder, retriever and a simplifier to simplify citations to improve citation quality. Another category related to citation generation is LLM attribution (Jain et al., 2023; Xu et al., 2024a; Gao et al., 2023a), which refers to the capacity of an LLM to generate and provide evidence (Li et al., 2023). For instance, Recitation Augmented Language Models(Sun et al., 2023), learns to sample documents from LLM’s self-knowledge and construct a path of attributing passages to generate the final answer.
@@
Need for comprehensive and fair comparisons: There is a lack of comprehensive and fair horizontal comparisons between various methods. Previous works have not been examined on new emerging Large Language Models, and since there are now some new metrics like citation granularity, a comprehensive evaluation is needed. Works following ALCE (Sun et al., 2024; Lee et al., 2024; Slobodkin et al., 2024) only compare their methods to ALCE baselines. Others (Asai et al., 2023; Sun et al., 2023) that focus more on answer accuracy are not evaluated on the citation benchmark, which means a lack of horizontal comparisons between SOTAs.
Given the problems above, a toolkit that unifies different methods is crucial for researchers to realize their workflow quickly and facilitate fair and straightforward horizontal comparisons between various methods and efficient improvement and innovation. However, there is still no such open-source toolkit to compare these state-of-the-art methods conveniently and fairly.
To this end, we present Citekit , an open-source, extensible, and user-friendly modular toolkit that can be easily used to construct a pipeline for citation generation by combining preset modules or propose an enhanced design by utilizing customized modules and modifying their interconnections accordingly.
@@
Citekit offers four different types of modules: Input , Generation Module , Enhancing Module , and Evaluator , which are combined in a pipeline. Input contains automatic components for loading data and making prompts, and is accessible by other modules. Generation Module is for response generation, where LLM follows the instructions and uses retrieved documents to generate an answer with citations or explicit reference to the documents for further process, featuring a wide range of LLM’s supported and adaptive generation modes, it can satisfy different need for various generation task. Enhancing Module contains some components that can assess, retrieve, plan, and edit, and they can be customized for different tasks and even combined into clusters. Evaluator integrates different metrics to evaluate the output of the pipeline and other new metrics could also be defined and utilized. For training-based methods, a parallel data export component can output evaluation results that could be an indicator of the performance of a method and can also be used as training data for supervised learning and Reinforcement Learning (RL). The toolkit also provides significant versatility in customizing new modules to quickly and conveniently realize an improved method. We provide 11 baseline recipes using Citekit to comprehensively and fairly compare these SOTA methods with the latest models Llama3 and GPT-4o across 8 various metrics on answer factuality and citation quality. Our experiments show that despite each method having its special strengths, it is challenging to achieve optimal results in both answer accuracy and citation quality. Therefore, we used Citekit to combine the most effective functional modules, achieving a balance in both answer accuracy and citation quality. Our contributions can be summarized as follows:
@@
  • We propose a framework that modularizes the citation tasks, with 4 main modules decomposing citation pipelines into request wrapping, generating, enhancing, and evaluating to unify different methods. The framework contains 14 components and 16 functions to define complicated interconnections between modules and satisfy different needs for citation generation tasks.
@@
  • We design and complete Citekit , an easy-to-use and extensible toolkit based on our framework to help reproduce, compare, and combine different methods. We pre-defined 11 recipes to cover 11 citation generation paradigms, all derived from state-of-the-art research.
@@
  • We conduct a comprehensive evaluation and comparison of the existing 11 baselines on 2 SOTA LLMs and propose an improved new method, self-RAG Snippet, by combining effective components. Our method achieves balance in both answer accuracy and citation quality, showing the convenience of verifying new methods.
@@ 4.1 Baselines and metrics
We evaluate 11 baselines in total using the state-of-the-art open-source and closed-source Large Language Models, GPT-4o (OpenAI, 2024) and Llama3-8B-Instruct (AI@Meta, 2024), respectively on ASQA dataset. ALCE Vanilla, Snippet, and Summ directly prompt the LLM to generate citations, using full documents, snippets, and summaries respectively. ALCE Interact (Gao et al., 2023b) uses document summaries and interactively provides full documents. AAR Lee et al. (2024) asked the LLM to revise the answer, while VTG Hennigen et al. (2024) will verify the answer and retrieve more supplementary documents for regeneration if needed. Citation Enhanced (Li et al., 2024b) method retrieves documents after generation, and Recitation Augmented (Sun et al., 2023) sample documents from self-knowledge of LLMs. Attribute First, then Generate (Slobodkin et al., 2024) will provide attributing spans to help the generation, while Blueprint Fierro et al. (2024) provides some questions to guide the generation. For self-RAG (Asai et al., 2023), we use our prompt version instead of a trained Model (Appendix A.2), to retrieve documents and generate sentence-by-sentence. To further improve this method on the citation granularity, we use snippets to make the model cite a smaller span as our improved method as self-RAG Snippet. We show the improved prompt in Appendix 13
As for metrics, we use metrics from ALCE, including fluency, correctness, rouge, citation recall, and precision. Besides, we also evaluate the correct prediction of citation need and the citation granularity.
@@ A.1 Blueprint Model
For the Blueprint Model, we use the abstractive model to produce general-purpose questions: the paragraph is the input and the question is the output. We use prompts to make LLMs generate questions. ALCE provides question-answer pairs for ASQA dataset, and in each pair the sub-question shows an aspect of answering the final question. We use these pairs to complete a 2-shot prompt for ICL. For answer generation, we adjust the ALCE prompt to make LLMs answer all the subquestions. We show our prompts in Figure 9, 10
@@ A.2 Prompt self-RAG
As for Llama3-8B and GPT-4o, there is no trained version for self-RAG, we use prompt to make the LLM retrieve documents and generate, then use an NLI model to evaluate if the document is supportive and the answer is useful, respectively in 3 segments. A reranker will find the best segment and the sentence is add to the answer. Similar to Attribute First, then Generate, We use generated sentences as prefix to complement the sentence-by-sentence iterative generation. We show our prompts in Figure 11, 12

### v2

@@ Abstract
The emerging paradigm of enabling Large Language Models (LLMs) to generate citations in Question-Answering (QA) tasks is lacking in a unified framework to standardize and fairly compare different citation generation methods, leading to difficulties in reproduction and evaluation. Therefore, we introduce Citekit, an open-source and modular toolkit designed to facilitate the implementation and evaluation of existing citation generation methods, while also fostering the development of new approaches to improve citation quality. This tool is highly extensible, allowing users to utilize 4 main modules and 14 components to construct a pipeline, evaluating an existing method or innovative designs. Our experiments with two state-of-the-art LLMs and 11 citation generation baselines demonstrate varying strengths of different modules in answer accuracy and citation quality improvement, as well as the challenge of enhancing granularity. Based on our analysis of the effectiveness of components, we propose a new method, PEEP, obtaining a balanced answer accuracy and citation quality. Citekit is released at https://github.com/SjJ1017/Citekit.
@@
Jiajun Shen1,3*, Tong Zhou1*, Yubo Chen1,2, Kang Liu1,2,4 1The Key Laboratory of Cognition and Decision Intelligence for Complex Systems Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3University of Chinese Academy of Sciences 4Shanghai Artificial Intelligence Laboratory shenjiajun21@mails.ucas.ac.cn, tong.zhou@ia.ac.cn {yubo.chen, kliu}@nlpr.ia.ac.cn
@@ 1 Introduction
Given the urgent need, ALCE (Gao et al., 2023b) developed some basic methods to enable LLMs to generate citations in QA tasks and propose metrics for evaluating the quality of citations. Following ALCE’s contribution, there are other methods that either use training (Huang et al., 2024a; Li et al., 2024a; Ye et al., 2024; Huang et al., 2024b) or construct complicated pipelines to enhance the ability of generative models in citing external documents (Zhang et al., 2024; Sun et al., 2024; Lee et al., 2023; Fierro et al., 2024; Qian et al., 2024). Another category related to citation generation is LLM attribution (Jain et al., 2023; Xu et al., 2024a; Gao et al., 2023a; Sun et al., 2023; Huang et al., 2024c; Cattan et al., 2024; Abolghasemi et al., 2024), which refers to the capacity of an LLM to generate and provide evidence (Li et al., 2023).
@@
Need for comprehensive and fair comparisons: There is a lack of comprehensive and fair horizontal comparisons between various methods. Previous works have not been examined on new LLMs, and since there are now some new metrics like citation granularity, a comprehensive evaluation is needed. Works following ALCE (Sun et al., 2024; Lee et al., 2024; Slobodkin et al., 2024) only compare their methods to ALCE baselines. Others (Asai et al., 2023; Sun et al., 2023) that focus more on answer accuracy are not evaluated on the citation benchmark, which means a lack of horizontal comparisons between SOTAs.
Given the problems above, a toolkit that unifies different methods is crucial for fast workflow implementation, fair comparisons between various methods, and efficient improvement and innovation. Therefore, we present Citekit , an open-source, extensible, and user-friendly modular toolkit to construct pipelines for citation generation.
@@
Citekit offers four different types of modules: Input , Generation Module , Enhancing Module , and Evaluator , which are combined in a pipeline. Input contains automatic components for loading data and making prompts, and is accessible by other modules. Generation Module is for response generation, where LLM follows the instructions and uses retrieved documents to generate an answer with citations or explicit reference to the documents for further process, featuring a wide range of LLM’s supported and adaptive generation modes, it can satisfy different need for various generation task. Enhancing Module contains some components that can assess, retrieve, plan, and edit, and they can be customized for different tasks and even combined into clusters. Evaluator integrates different metrics to evaluate the output of the pipeline and other new metrics could also be defined and utilized. For training-based methods, a parallel data export component can output evaluation results for supervised learning and Reinforcement Learning (RL). Citekit also provides significant versatility in customizing new modules to quickly and conveniently realize an improved method. We provide 11 baseline recipes using Citekit to comprehensively and fairly compare these SOTA methods with the latest models Llama3 and GPT-4o, showing the strength of different modules and remaining challenges. Finally, we learn from the baselines to combine the most efficient planning module, reviser, and simplifier to build our method, PEEP. We achieve a balance in both answer accuracy and citation quality. Our contributions can be summarized as follows:
@@
  • We propose a framework that modularizes citation tasks, with four main modules decomposing citation pipelines into request wrapping, generating, enhancing, and evaluating to unify different methods. The framework contains 14 components and 16 functions to define complicated interconnections between modules and satisfy different needs for citation generation tasks.
@@
  • We design and complete Citekit , an easy-to-use and extensible toolkit based on our framework to help reproduce, compare, and combine different methods. We pre-defined 11 recipes to cover 11 citation generation paradigms, all derived from SOTA research.
@@
  • We conduct a comprehensive evaluation and comparison of the existing 11 baselines on 2 SOTA LLMs and propose an improved new method, PEEP, by combining effective components. Our method achieves balance in answer accuracy and citation quality, showing the convenience of verifying new methods.
@@ 4.1 Baselines and metrics
We evaluate 11 baselines in total using the state-of-the-art open-source and closed-source LLMs, GPT-4o (OpenAI, 2024) and Llama3-8B-Instruct (AI@Meta, 2024) on ASQA dataset. ALCE Vanilla, Snippet, and Summ directly prompt the LLM to generate citations using full documents, snippets, and summaries respectively. ALCE Interact (Gao et al., 2023b) uses document summaries and interactively provides full documents. AAR Lee et al. (2024) asked the LLM to revise the answer, while VTG Sun et al. (2024) will verify the answer and retrieve more supplementary documents for regeneration. Citation Enhanced (Li et al., 2024b) method retrieves documents after generation, and Recitation Augmented (Sun et al., 2023) sample documents from pre-training data. Attribute First, then Generate (Slobodkin et al., 2024) and Blueprint Fierro et al. (2024) provides some attributing spans or questions to guide the generation. For self-RAG (Asai et al., 2023), we use our prompt version instead of a trained model to retrieve documents and generate sentence-by-sentence. We use metrics from ALCE for evaluation, including fluency, correctness, rouge, citation recall, and precision. We also evaluate the appropriate citation rate and the citation granularity.
4.2 Our method
We build our new pipeline, PEEP, combining the most efficient modules. We use a Planner to decompose the question in two as much as atomic questions (Yan et al., 2024), and two Editors (reviser and simplifier) to combine the answers for each atomic question and simplify the citation afterward. We use Parallel generation to get answers with sufficient diversity, as shown in Figure 5.
Figure 5: Design of PEEP. We show an example of generating a comprehensive answer for an ASQA question.
@@ A.1 Blueprint Model
For the Blueprint Model, we use the abstractive model to produce general-purpose questions: the paragraph is the input and the question is the output. We use prompts to make LLMs generate questions. ALCE provides question-answer pairs for ASQA dataset, and in each pair the sub-question shows an aspect of answering the final question. We use these pairs to complete a 2-shot prompt for ICL. For answer generation, we adjust the ALCE prompt to make LLMs answer all the subquestions. We show our prompts in Figure 10, 11
@@ A.2 Prompt self-RAG
As for Llama3-8B and GPT-4o, there is no trained version for self-RAG, we use prompt to make the LLM retrieve documents and generate, then use an NLI model to evaluate if the document is supportive and the answer is useful, respectively in 3 segments. A reranker will find the best segment and the sentence is add to the answer. Similar to Attribute First, then Generate, We use generated sentences as prefix to complement the sentence-by-sentence iterative generation. We show our prompts in Figure 12, 13